OpenStudy (anonymous):

2(3x - 7) + 4 (3 x + 2) = 6 (5 x + 9 ) + 3

OpenStudy (anonymous):

Which part are you stuck on I will try to help

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

Hmmmm.. ok are you familiar with pemdas?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

got to go ill do this later but thanks for the help

OpenStudy (anonymous):

2(3x - 7) + 4(3x + 2) = 6(5x + 9) + 3 6x - 14 + 12x + 8 = 30x + 54 + 3 {used distributive property} 18x - 6 = 30x + 57 {combined like terms} -6 = 12x + 57 {subtracted 18x from both sides} -63 = 12x {subtracted 57 from both sides} x = -63/12 {divided both sides by 12} x = -21/4 {reduced}

OpenStudy (anonymous):

x = -5.25

OpenStudy (anonymous):

Color me impressed @AskMeLotss

OpenStudy (anonymous):

XD thanks